# Look as You Leap: Planning Simultaneous Motion and Perception for High-d o f Robots

### v1

@@ Abstract
In this work, we address the problem of planning robot motions for a high-degree-of-freedom (d o f) robot that effectively achieves a given perception task while the robot and the perception target move in a dynamic environment. Achieving navigation and perception tasks simultaneously is challenging, as these objectives often impose conflicting requirements. Existing methods that compute motion under perception constraints fail to account for obstacles, are designed for low-d o f robots, or rely on simplified models of perception. Furthermore, in dynamic real-world environments, robots must replan and react quickly to changes and directly evaluating the quality of perception (e.g., object detection confidence) is often expensive or infeasible at runtime. This problem is especially important in human-centered environments such as homes and hospitals, where effective perception is essential for safe and reliable operation. To address these challenges, we propose a GPU-parallelized perception-score-guided probabilistic roadmap planner with a neural surrogate model (ps-prm). The planner explicitly incorporates the estimated quality of a perception task into motion planning for high-d o f robots. Our method uses a learned model to approximate perception scores and leverages GPU parallelism to enable efficient online replanning in dynamic settings. We demonstrate that our planner, evaluated on high-d o f robots, outperforms baseline methods in both static and dynamic environments in both simulation and real-robot experiments.
@@ I Introduction
Achieving mobile manipulation (e.g., navigating to a table to grasp a tool) alongside perception tasks (e.g., tracking an object or a human) remains highly challenging. In this work, we address the problem of planning robot motions for a high-degree-of-freedom (d o f) robot that effectively achieves a given perception task while the robot and the perception target move in a dynamic environment. A fundamental challenge emerges: the objectives of motion and perception often conflict—perception tasks impose non-trivial constraints on robot motion, as performance varies with object shape, occlusion, distance, and viewing angle [1, 2, 3]. For example, detecting human faces often requires maintaining a frontal view, while reliable object tracking typically benefits from close proximity. At the same time, motion planning for high-degree-of-freedom (d o f) robots must follow kinematic constraints, avoid obstacles and nearby humans, and account for the complex, nonlinear relationship between the robot’s configuration and the camera pose.
@@
As an example, consider the situation in Fig. 1. A mobile robot with wrist-mounted camera must navigate a cluttered environment while observing a monitor at the end of the hallway. The robot must navigate through clutter with a differential drive base while simultaneously maintaining tracking performance, i.e., it avoids occlusions and attains high tracking confidence.
@@
Figure 1: The robot—a UR5 mounted on a differential drive base—navigates a cluttered environment while maintaining high perception performance. In this example, the robot must continuously track a monitor located down the perpendicular hallway using its wrist-mounted camera. Top: Illustration of the robot’s trajectory—fading from right to left—as it moves while maintaining a high detection rate of the monitor. Bottom: Camera views corresponding to each robot configuration shown in the top figure.
@@
These challenges are further compounded in dynamic environments, where objects and people move. In such settings, robots must continuously adapt to changes while preserving high perception quality [4, 5, 6, 7]. For example, a key motivation of this work comes from human-centered environments such as nurse training settings [8, 9], where robotic tutors are used to support skill acquisition. In such a system [10], the robot must continuously monitor a trainee to detect compliance with sterile techniques and provide timely feedback while simultaneously navigating the environment. Fig. 7 and Fig. 10 illustrate this challenge, where the robot must continuously track nurses’ faces while navigating through a cluttered clinical environment. However, directly evaluating perception performance (e.g., detection confidence) in such settings can be prohibitively expensive or even impossible. This motivates the need for efficient mechanisms to approximate perception quality and support fast, responsive replanning in high-dimensional spaces.
@@
How do existing planning methods tackle planning for perception in high d o f setting? Existing planning methods that consider both motion and perception have several limitations. Informative path planners [11, 12, 13, 14] aim to find paths that maximize information gain while respecting constraints like path length or budget, but they often become computationally expensive, especially when the problem size grows. Perception-aware path planners are typically designed for low-d o f systems [15, 16, 17, 18, 19, 20] or rely on overly simplified perception models [21], such as keeping a point of interest centered in the field of view. Active perception approaches [22, 23, 24, 25, 26, 27, 28, 29, 30] focus on selecting discrete viewpoints to maximize scene understanding, rather than ensuring consistent perception quality along a motion trajectory. Moreover, most of these approaches are not designed to operate in dynamic environments where objects or humans may move and where continuous adaptation is needed. As a result, they lack the responsiveness required for real-time perception-aware planning in practical settings.
@@
In this paper, we propose Perception-Score-guided Probabilistic Roadmap Planning (PS-PRM)—a GPU-parallelized, roadmap-based planner that explicitly incorporates the estimated quality of a given perception task (e.g., object detection confidence) into motion planning for high-degree-of-freedom (DOF) robots. The perception score is predicted by a neural surrogate model trained to approximate perception scores for specific tasks. ps-prm jointly optimizes for both motion and perception by associating each configuration with an estimated perception score and using these scores to guide roadmap construction and path selection. To efficiently approximate perception quality during planning, we introduce a neural surrogate model trained on data from various objects and human targets. To account for occlusions in cluttered environments, we incorporate a ray-casting-based pipeline that dynamically adjusts perception estimates. Finally, to enable fast replanning in dynamic environments, we develop a GPU-parallelized framework that executes collision checking, forward kinematics (FK), and perception evaluation in batch. These components allow our method to scale to high-dimensional planning problems while maintaining robust perception performance in both static and dynamic scenarios.
@@
  • Extensive simulation and real-robot experiments on three different high-d o f robots, demonstrating that ps-prm consistently improves both perception performance and planning efficiency compared to baseline methods.
@@ II Problem Statement
We consider a robot with $k$ controllable joints, such as a mobile robot with an onboard camera with limited field of view. The robot is described by its configuration space $\mbox{$\mathcal{Q}$}\subseteq\mathbb{R}^{k}$. The robot is located in a workspace $\mbox{$\mathcal{W}$}\subseteq\mathbb{R}^{3}$ that contains objects and people of interest that may need to be monitored, as well as obstacles to be avoided. The robot has access to the pose and geometry of all obstacles in $\mathcal{W}$ and the positions of the centroids of objects of interest. As the robot moves, the objects or humans of interest may also move, and we assume the robot has real-time access to their poses. The subset of all robot configurations that are collision-free is denoted by $\mbox{$\mathcal{Q}_{free}$}\subseteq\mbox{$\mathcal{Q}$}$. The motion planning problem is to find a path $\sigma:[0,1]\rightarrow\mbox{$\mathcal{Q}_{free}$}$ that assigns a collision-free configuration to each input parameter, starting from $\sigma(0)=\mbox{$q_{start}$}$ and ending in a goal region $\sigma(1)\in\mbox{$\mathcal{Q}_{goal}$}$.
We denote the set of all paths in $\mathcal{Q}$ as $\Sigma_{\mathcal{Q}}$, and all collision-free paths as $\Sigma_{\mathcal{Q}_{\mathrm{free}}}$. We additionally define a motion cost function $c_{m}:\Sigma_{\mathcal{Q}}\rightarrow\mathbb{R}^{+}$ and a perception cost function $c_{p}:\Sigma_{\mathcal{Q}}\rightarrow\mathbb{R}^{+}$. The objective of this work is to find a collision-free path $\sigma^{*}:[0,1]\rightarrow\mathcal{Q}_{\text{free}}$, given a path planning problem $(\mathcal{Q}_{\text{free}},q_{\text{start}},q_{\text{goal}})$, that minimizes a composite cost function $c(\sigma)=c_{m}(\sigma)+\alpha\cdot c_{p}(\sigma)$, where the trade-off between motion and perception cost is governed by a scalar weight $\alpha\geq 0$. The optimal path is given by $\sigma^{*}=\arg\min_{\pi\in\Sigma_{\mathcal{Q}_{\text{free}}}}\left[c_{m}(\sigma)+\alpha\cdot c_{p}(\sigma)\right]$, if such a path exists.
@@ III-A Informative Path Planning
Another related line of works is the planning under uncertainty or planning in information spaces, which focuses on selection of trajectories that minimize the localization uncertainty. This problem has generally been solved with Partially Observable Markov Decision Processes (POMDPs) or through graph-search in the belief space [38]. While these approaches are well-established, in general their computational complexity grows exponentially in the number of possible actions and observations.
To overcome these issues, the use of sampling-based motion planning algorithms, such as the RRT [39] and the probabilistic roadmap (PRM) [40], has increased enormously in past years. Such algorithms have the advantage of quickly exploring a space of interest to achieve a feasible solution. Variants of sampling-based motion planning have also been proposed to maximize some metric of information quality. The prior algorithms have been shown to solve related problems for certain objective functions, including information-rich RRTs [41] using Fisher information matrices and belief-space RRTs [42]. While information-rich sampling based motion planners mentioned above like belief-space RRTs are well-suited for energy minimization tasks, in this work, we specifically focus on selecting trajectories that maximize the perception task performance without considering robot control efforts. Furthermore, these prior algorithms often focus on 2D environments [41, 42, 32, 33] and their information metrics do not apply to the perception task we consider in this work where the approximate analytical model suffers.
@@ III-B Perception-aware Path Planning
For UAVs, most methods formulate the problem as optimization using Model Predictive Control (MPC) [15, 44, 11, 14, 19] or Reinforcement Learning (RL) [12, 17, 18, 49]. MPC-based methods jointly optimize perception and motion for agile UAV navigation using numerical optimization, without obstacles [15] or for static obstacles and occlusions [44]. For example, the work [21] steer the UAV towards informative areas using a path planner while keeping landmarks in the camera field of view through optimization. RL-based methods create policies that directly map from sensor measurements to controls while keeping visibility of regions of interest. The policy learner has access to privileged information during training, e.g., by using MPC [50], a planner [17] or a state-based trained expert [18]. Another line of work focuses on using the UVA or MVA for autonomous aerial cinematography or low altitude traffic surveillance that can reason about both geometry and scene context in real-time [51, 52].
Unlike these works, we consider high-d o f systems—e.g., mobile robots with an onboard camera with limited field-of-view controlled by multiple joints—with kinematic constraints on environments with many obstacles, where optimization is likely to get stuck in a local minima. We focus on simultaneous motion and perception high-d o f robots such as mobile manipulators instead of UAVs or MAVs. Furthermore, unlike most of the perception-aware planning works that aim at maintaining visibility of landmarks and informative regions for visual odometry, we address more complex perception tasks which are significantly more computationally expensive to evaluate.
@@ III-C Active Perception
Similar to our work, active perception considers motion costs and quality of the perception task. However, the main focus of next best view planning is proposing and selecting views from which a scene can be efficiently observed while in this paper we are interested in computing trajectories from a start towards a predefined goal while striking a good balance between the robot perception performance and most cost. In addition, one of the most important goals of active perception is active localization, which seeks to compute control actions and trajectories that minimize the pose estimation uncertainty. In contrast, our work focuses more on planning for simultaneous perception and motion in a goal-reaching task.
@@ IV Overview of PS-PRM
Figure 4: Pipeline for perception-aware evaluation using a neural surrogate model and occlusion-aware ray tracing. From left to right: the environment contains the target and an occluding object; a surrogate model predicts a dense perception score field; ray tracing incorporates occlusion to refine the score; and camera views at viewpoints A and B illustrate visibility differences. High perception scores correspond to views with minimal occlusion.

### v2

@@ Abstract
Most common tasks for robots in dynamic spaces require that the environment is regularly and actively perceived, with many of them explicitly requiring objects or persons to be within view, i.e., for monitoring or safety. However, solving motion and perception tasks simultaneously is challenging, as these objectives often impose conflicting requirements. Furthermore, while robots must react quickly to changes in the environment, directly evaluating the quality of perception (e.g., object detection confidence) is often expensive or infeasible at runtime. This problem is especially important in human-centered environments, such as homes and hospitals, where effective perception is essential for safe and reliable operation. In this work, we address the challenge of solving motion planning problems for high-degree-of-freedom (d o f) robots from a start to a goal configuration with continuous perception constraints under both static and dynamic environments. We propose a GPU-parallelized perception-score-guided probabilistic roadmap planner with a neural surrogate model (ps-prm). Unlike existing active perception-, visibility-aware or learning-based planners, our work integrates perception tasks and constraints directly into the motion planning formulation. Our method uses a neural surrogate model to approximate perception scores, incorporates them into the roadmap, and leverages GPU parallelism to enable efficient online replanning in dynamic settings. We demonstrate that our planner, evaluated on high-d o f robots, outperforms baseline methods in both static and dynamic environments in both simulation and real-robot experiments.
@@ I Introduction
Achieving mobile manipulation (e.g., navigating to a table to grasp a tool) alongside perception tasks (e.g., tracking an object or a human) remains highly challenging. In this work, we address the problem of solving motion planning problems for high degree-of-freedom (d o f) robots from a start to a goal configuration with a given perception task under both static and dynamic environments. A fundamental challenge emerges: the objectives of motion and perception often conflict. The perception tasks could impose non-trivial constraints on robot motion, as performance varies with object shape, occlusion, distance, and viewing angle [104, 107, 58]. For example, achieving reliable perception may require the robot to move closer to a target or adopt specific viewing angles, which can conflict with motion objectives such as maintaining clearance from obstacles, respecting kinematic limits, or following a short and smooth path. At the same time, planning for high-degree-of-freedom (d o f) robots with perception constraints must account for the complex, nonlinear relationship between the robot’s configuration and the camera pose.
@@
As an example, consider the situation in Fig. 1. A mobile manipulator with a wrist-mounted camera must move from a start to a goal configuration through a cluttered environment while observing a monitor at the end of the hallway. The robot must reach its goal with a trajectory that maintains good tracking performance, i.e., it avoids occlusions and achieves high tracking confidence.
@@
Figure 1: The robot (a UR5 mounted on a differential drive base) moves through a cluttered environment while maintaining high perception performance. In this example, the robot must continuously track a monitor located down the perpendicular hallway using its wrist-mounted camera. Top: Illustration of the robot’s trajectory (fading from right to left) as it moves while maintaining a high detection rate of the monitor. Bottom: Camera views corresponding to the robot configurations shown in the top figure.
@@
The planning challenges are further compounded in dynamic environments, where objects and people move. In such settings, robots must continuously adapt to changes while preserving high perception quality [62, 10, 69, 56]. For example, a key motivation of this work comes from human-centered environments such as nurse training settings [4, 71], where robotic tutors are used to support skill acquisition. In such a system [70], the robot must continuously monitor a trainee to detect compliance with sterile techniques and provide timely feedback while simultaneously moving through the environment. Fig. 7 and Fig. 11 illustrate this challenge, where the robot must continuously track nurses’ faces while navigating through a cluttered clinical environment. However, naively evaluating perception performance (e.g., detection confidence) for perception-aware planning is prohibitively expensive in real environments [83] because it would require physically placing the sensor at a large number of distinct poses throughout the environment. This motivates the need for efficient mechanisms to approximate perception quality and support fast, responsive replanning in high-dimensional spaces.
@@
Our contribution focuses on formulating and addressing a new motion–perception planning problem: motion planning for a high-d o f robot with continuous perception constraints from a start to a goal configuration. Existing works that consider both motion and perception are not designed for the problem examined in this paper. Informative path planners [92, 84, 54, 53] aim to find paths that maximize information gain while respecting constraints such as path length or budget, but they often become computationally expensive, especially when the problem size grows. Perception-aware path planners are typically designed for low-d o f systems [25, 33, 50, 86, 61, 108] or rely on overly simplified perception models [8], such as keeping a point of interest centered in the field of view. Active perception approaches [7, 3, 85, 48, 20, 94, 60, 77, 59] focus on selecting discrete viewpoints to maximize scene understanding, rather than ensuring consistent perception quality along a motion trajectory. Moreover, most of these approaches are not designed to operate in dynamic environments where objects or humans may move and where continuous adaptation is needed. As a result, they lack the responsiveness required for real-time perception-aware planning in practical settings.
@@
In this paper, we propose Perception-Score-guided Probabilistic Roadmap Planning (PS-PRM): a GPU-parallelized, roadmap-based planner that explicitly incorporates the estimated quality of a given perception task (e.g., object detection confidence) into motion planning for high-degree-of-freedom (DOF) robots. The perception score is predicted by a neural surrogate model trained to approximate perception scores for specific tasks. A key aspect of our formulation is that perception quality is modeled in SE(3) space rather than in robot-specific configuration space. By basing the surrogate perception model on the camera pose and its relative transform to the target, it becomes robot- and environment-agnostic, enabling direct reuse across different high-d o f platforms with similar sensor setups. Fast forward kinematics (FK) allow for this model to still be directly and efficiently linked to the configuration space. This design choice allows our planner to generalize beyond a single robot and environment, a property we demonstrate in our experiments. ps-prm jointly optimizes for both motion and perception by associating each configuration with an estimated perception score and using these scores to guide roadmap construction and path selection. To efficiently approximate perception quality during planning, our neural surrogate model is trained on data from various objects and human targets that are typically present in the specific environment the robot is expected to operate. To account for occlusions in cluttered environments, we incorporate a ray-casting-based pipeline that dynamically adjusts perception estimates. Finally, to enable fast replanning in dynamic environments, we develop a GPU-parallelized framework that executes collision checking, FK, and perception evaluation in batch. These components allow our method to scale to high-dimensional planning problems while maintaining robust perception performance in both static and dynamic scenarios. Existing visibility-aware, informative, or active-perception planners do not incorporate perception score surrogates with a direct link to the robot’s configuration space, nor do they support real-time replanning for high-d o f robots in dynamic environments. In contrast, our framework explicitly couples a perception-score surrogate, occlusion-aware ray casting, and roadmap-based planning, all accerelated with GPU parallelization.
@@
  • Extensive simulation and real-robot experiments on four different high-d o f robot systems, demonstrating that ps-prm consistently improves both perception performance and planning efficiency compared to baseline methods.
@@ II Problem Statement
We consider a robot with $k$ controllable joints, such as a mobile robot with an onboard camera with limited field of view. The robot’s configuration space is $\mbox{$\mathcal{Q}$}\subseteq\mathbb{R}^{k}$. The robot is located in a workspace $\mbox{$\mathcal{W}$}\subseteq\mathbb{R}^{3}$ that contains objects and people of interest that may need to be monitored, as well as obstacles to be avoided. The robot has access to the pose and geometry of all obstacles in $\mathcal{W}$ and the positions of the centroids of objects of interest. As the robot moves, the objects or humans of interest may also move, and we assume the robot has real-time access to their poses. The subset of all robot configurations that are collision-free is denoted by $\mbox{$\mathcal{Q}_{free}$}\subseteq\mbox{$\mathcal{Q}$}$. The motion planning problem is to find a path $\sigma:[0,1]\rightarrow\mbox{$\mathcal{Q}_{free}$}$ that assigns a collision-free configuration to each input parameter, starting from $\sigma(0)=\mbox{$q_{start}$}$ and ending in a goal region $\sigma(1)\in\mbox{$\mathcal{Q}_{goal}$}$.
We denote the set of all paths in $\mathcal{Q}$ as $\Sigma_{\mathcal{Q}}$, and all collision-free paths as $\Sigma_{\mathcal{Q}_{\mathrm{free}}}$. We additionally define a motion cost function $c_{m}:\Sigma_{\mathcal{Q}}\rightarrow\mathbb{R}^{+}$ and a perception cost function $c_{p}:\Sigma_{\mathcal{Q}}\rightarrow\mathbb{R}^{+}$. The problem we solve is to find a collision-free path $\sigma^{*}:[0,1]\rightarrow\mathcal{Q}_{\text{free}}$, given a path planning problem $(\mathcal{Q}_{\text{free}},q_{\text{start}},q_{\text{goal}})$, that minimizes a composite cost function $c(\sigma)=c_{m}(\sigma)+\alpha\cdot c_{p}(\sigma)$, where the trade-off between motion and perception cost is governed by a scalar weight $\alpha\geq 0$. The optimal path is given by $\sigma^{*}=\arg\min_{\pi\in\Sigma_{\mathcal{Q}_{\text{free}}}}\left[c_{m}(\sigma)+\alpha\cdot c_{p}(\sigma)\right]$, if such a path exists.
@@ III Related Work
In this section, we discuss three categories of related problems examined in this paper: informative path planning, perception-aware path planning, and active perception. While all of these approaches consider perception and motion, they address problem formulations that are fundamentally different from ours.
@@ III-A Informative Path Planning
Another related line of work is planning under uncertainty or planning in information spaces, which focuses on selection of trajectories that minimize the localization uncertainty. This problem has generally been solved with Partially Observable Markov Decision Processes (POMDPs) or through graph-search in the belief space [12]. While these approaches are well-established, in general their computational complexity grows exponentially in the number of possible actions and observations.
To overcome these issues, the use of sampling-based motion planning algorithms, such as the RRT [43] and the probabilistic roadmap (PRM) [36], has increased enormously in past years. Such algorithms have the advantage of quickly exploring a space of interest to achieve a feasible solution. Variants of sampling-based motion planning have also been proposed to maximize some metric of information quality. The prior algorithms have been shown to solve related problems for certain objective functions, including information-rich RRTs [44] using Fisher information matrices and belief-space RRTs [15].
Unlike these works, this paper specifically focuses on selecting a trajectory that balances the perception task performance with motion cost. Furthermore, these prior algorithms often focus on 2D environments [44, 15, 30, 75] and their information metrics do not apply to the perception task we consider in this work where the approximate analytical model is infeasible.
@@ III-B Perception-aware Path Planning
For UAVs, most methods formulate the problem as optimization using Model Predictive Control (MPC) [25, 66, 92, 53, 61] or Reinforcement Learning (RL) [84, 50, 86, 93]. MPC-based methods jointly optimize perception and motion for agile UAV navigation using numerical optimization, without obstacles [25] or for static obstacles and occlusions [66]. For example, the work [8] steer the UAV towards informative areas using a path planner while keeping landmarks in the camera field of view through optimization. RL-based methods create policies that directly map from sensor measurements to controls while keeping visibility of regions of interest. The policy learner has access to privileged information during training, e.g., by using MPC [106], a planner [50] or a state-based trained expert [86]. Another line of work focuses on using the UAV or MAV for autonomous aerial cinematography or low altitude traffic surveillance that can reason about both geometry and scene context in real-time [11, 14].
Unlike these works, we consider high-d o f systems (e.g., mobile robots with an onboard camera with limited field-of-view controlled by multiple joints) with kinematic constraints on environments with many obstacles, where optimization is likely to get stuck in a local minima. We focus on simultaneous motion and perception high-d o f robots such as mobile manipulators instead of UAVs or MAVs. Furthermore, unlike most of the perception-aware planning works that aim at maintaining visibility of landmarks and informative regions for visual odometry, we address more complex perception tasks which are significantly more computationally expensive to evaluate.
Method Fixed Goal Cont. Perc. General Perc. High-DoF
Ours
Active Perception
Perception-aware Planning
TABLE I: Comparison between our perception-constrained motion planner, active perception, and perception-aware path planning. Fixed Goal indicates whether the task specifies a fixed goal configuration. Cont. Perc. (continuous perception) denotes that perception constraints are enforced throughout the entire trajectory rather than only at discrete waypoints. General Perc. (general perception) indicates support for general perception objectives beyond a single task category (e.g., not limited to mapping or exploration). High-DoF denotes applicability to high-degree-of-freedom robotic systems.
@@ III-C Active Perception
Recently, learning-based active perception [90, 96, 34, 17, 101] has become popular across robotics. These approaches primarily formulate perception as an information-gathering problem, finding viewpoints that reduce uncertainty in mapping, object state estimation, or task-relevant observations. For example, [90] formulates damage inspection as a POMDP and uses deep reinforcement learning to select next sensing poses that maximize crack-segmentation quality during underwater structural inspection. The authors in [96] propose an asynchronous active vision–action framework that serially trains a next-best-view camera policy and a next-best-pose gripper policy for manipulation tasks under occlusion, but the perception policy is still defined in a low-dimensional camera-space rather than full robot configuration space. Several mapping-driven approaches [34, 17] focus on mobile robots or aerial platforms, where active perception guides viewpoint selection to improve scene reconstruction fidelity. [34] uses confidence-driven next-best-views to actively densify Gaussian-splatting maps, while [17] develops a real-time navigation pipeline in the maps that couples safe corridor planning with fast vision-based localization. Broader surveys such as [101] further emphasize how vision-based learning in drones centers around improving perception for navigation, mapping, and control, typically in 3-d o f mobile systems.
Unlike learning-based active perception methods which choose perception-oriented intermediate goals or viewpoints, our planner operates with a fixed motion planning goal. In addition, active perception is largely decoupled from whole-body motion planning and focuses on viewpoint selection for mapping or manipulation in low-dimensional action spaces. This is different from our objective of embedding perception constraints directly within high-d o f motion planning, which requires sensing quality, occlusion reasoning, and task performance to be jointly evaluated in the robot’s full configuration space. In Table I, we summarize the different objectives of our method, active perception and perception-aware path planning.
@@ IV Overview of PS-PRM
Figure 4: Pipeline for perception-aware evaluation using a neural surrogate model and occlusion-aware ray tracing. From left to right: the environment contains the target (human) and an occluding object (orange box); a surrogate model predicts a dense perception score field; ray tracing incorporates occlusion to refine the score; and camera views at viewpoints A and B illustrate visibility differences. High perception scores correspond to views with minimal occlusion.